OpenStudy (anonymous):

Find the rate at which the surface area of a sphere is changing with respect to its radius r when r = 3 cm.

OpenStudy (radar):

What is the rate of the radius change (dr/dt)?

OpenStudy (radar):

If R is a constant 3, then the surface area would also be a constant value.

OpenStudy (anonymous):

\[ A= 4 \pi r^2\\ \frac {dA}{dt}= 8 \pi r \frac {dr}{dt}=24 \pi \frac {dr}{dt}\\ \] If you know \( \frac{dr}{dt}\), then you would know \( \frac {dA}{dt}\)

OpenStudy (anonymous):

when r=3

OpenStudy (anonymous):

Perhaps all that was sought was the formula for dA/dr.

OpenStudy (anonymous):

It could be. The question must have been more precise. @douglaswinslowcooper

OpenStudy (radar):

That certainly answered the question. I thought there was more to it. But after re-reading that was all that was needed. well done @eliassaab